OpenStudy (anonymous):

3. The diagonals of an isosceles trapezoid ,BD and AC, are labeled 5x and 3x + 40. Find the length of the diagonals. 3. The diagonals of an isosceles trapezoid ,BD and AC, are labeled 5x and 3x + 40. Find the length of the diagonals. @Mathematics

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how did you get 5

OpenStudy (anonymous):

becuz i am awesome

OpenStudy (anonymous):

.........

OpenStudy (anonymous):

can you please show me how to do it...?

OpenStudy (anonymous):

do not second guess me mortal

OpenStudy (anonymous):

I guess it to be 20

OpenStudy (anonymous):

............ in any other case i would come back at you with disgust for mortals but i really have to get this done so can you please help me

OpenStudy (anonymous):

ok i shall help you

OpenStudy (anonymous):

i think my work here is done

OpenStudy (anonymous):

o.O dear shabeer please help

OpenStudy (anonymous):

diagonals of an isosceles trapezoid are equal in length

OpenStudy (anonymous):

As this is an isosceles trapezoid the line segments AD and BC are equal in length. So I suppose Diagonal to be equal and 5x = 3x+40. A linear equation in one variable.

OpenStudy (anonymous):

5x and 3x + 40. so 5x = 3x + 40 5x - 3x = 40 x = 40/2 x = 20

OpenStudy (anonymous):

so 5x = 5*20 = 100 3x + 40 = 3*20 + 40 = 60 + 40 = 100

OpenStudy (anonymous):

Your right, Harikat. I too wanted to type the same thing.

OpenStudy (anonymous):

Ohh.... we were working on parallel tracks....☺

OpenStudy (anonymous):

lozl the both of u r so cute haha

OpenStudy (anonymous):

In your profile : you said you hate math. It's such an excellent subject. It's my favourite

OpenStudy (smartbrain):

@Atsie this is a question I am stumbled on.

OpenStudy (smartbrain):

In this step 5x and 3x + 40. so 5x = 3x + 40 5x - 3x = 40 In this step why did they switch the equals sign and how do you know its minus not plus. @Atsie x = 40/2 x = 20

Atsie (atsie):

@smartbrain I gotta go right now, I'll be back in an hour and then I'll answer ok?

OpenStudy (smartbrain):

I will probably be gone, but sure, that will help explain it. Ill see if somebody else can explain it meanwhile.

OpenStudy (smartbrain):

@AihberKhan can you possibly explain?

OpenStudy (aihberkhan):

@smartbrain Okay, so they are simply subtracting \(3x\) from both sides. Since they are subtracting, the positive \(3x\) becomes negative. You see? Also, they did not switch the equals sign they subtracted \(3x\) from both sides.

OpenStudy (smartbrain):

How did they conclude that the 3x is to be subtracted from both sides and not the 5x?

OpenStudy (aihberkhan):

If the \(5x\) is subtracted from both sides, then there will only be ONE side! Because if the \(5x\) was subtracted it would be \(=3x - 40 - 5x\) which simply can't happen.

OpenStudy (smartbrain):

I see, so you are trying to eliminate the x in the problem? Right?

OpenStudy (aihberkhan):

You are not trying to eliminate it. You are trying to isolate it. :) @smartbrain

OpenStudy (smartbrain):

Ok this answers my question. Thank you.

OpenStudy (aihberkhan):

No problem! :) Hope this helped! Have a great day! :) Also a medal would be much appreciated! Just click best response next to my answer. Thank You! Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @smartbrain